# A new intersection condition in extremal set theory

### v1

@@ Abstract.
We call a family $\mathcal{F}$ $(3,2,\ell)$-intersecting if $|A\cap B|+|B\cap C|+|C\cap A|\geq\ell$ for all $A$, $B$, $C\in\mathcal{F}$. We try to look for the maximum size of such a family $\mathcal{F}$ in case when ($\mathcal{F}\subset{[n]\choose k}$ or $\mathcal{F}\subset 2^{[n]}$. In the uniform case we show that if $\mathcal{F}$ is $(3,2,2)$-intersecting, then ||($|\mathcal{F}|\leq{n+1\choose k-1}+{n\choose k-2}$ and if $\mathcal{F}$ is $(3,2,3)$-intersecting, then ||($|\mathcal{F}|\leq{n\choose k-1}+2{n\choose k-3}+3{n-1\choose k-3}$. For the lower bound we construct a $(3,2,\ell)$-intersecting family and we show that this bound is sharp when $\ell=2$ or $3$ and $n$ is sufficiently large compared to $k$. In the non-uniform case we give an upper bound for a $(3,2,n-x)$-intersecting family, when $n$ is sufficiently large compared to $x$.
@@ 1. Introduction
[2] Let $k$, $n\geq 2k$ be integers and (${\mathcal{F}}\subset{[n]\choose k}$ be an intersecting family. Then $|{\mathcal{F}}|\leq\displaystyle\binom{n-1}{k-1}$.
@@
It was proved by Erdős, Ko, and Rado in their seminal paper [2] that for any $1\leq t<k$, the largest $t$-intersecting families ($\mathcal{F}\subset{[n]\choose k}$ are the maximal trivially $t$-intersecting ones if $n$ is large enough. The threshold function $n_{0}(k,t)$ of this property was determined by Wilson [18] after Frankl [5] settled the cases $t\geq 15$:
@@
The non-uniform version of the theorem for $t=2$ was proved by Katona.
@@
||$|\mathcal{F}|\leq\left\{\begin{array}[]{c c}\sum_{i=\frac{n+t}{2}}^{n}{n\choose i}&\text{if $n+t$ is even}\vspace{1mm}\\ \vspace{1mm}\sum_{i=\frac{n+t+1}{2}}^{n}{n\choose i}+{n-1\choose\frac{n+t-1}{2}}&\text{if $n+t$ is odd}\end{array}\right.$
@@
[3] $m(n,k,s)=\max\left\{{n\choose k}-{n-s\choose k},{k(s+1)-1\choose k}\right\}$
It was one of the favourite problems of Erdős and there was hardly a combinatorial lecture of him where he did not mention it. The Erdős Matching Conjecture, or EMC for short, is trivial for $k=1$ and was proved by Erdős and Gallai [4] for $k=2$ and it was settled in the case $k=3$ and $n\geq 4s$ in [12], for $k=3$, all $n$ and $s\geq s_{0}$ in [17], and finally, it was completely resolved for $k=3$ in [8].
The case $s=1$ is the classical Erdős-Ko-Rado theorem. In his original paper, Erdős proved it for $n\geq n_{0}(k,s)$ for some $n_{0}(k,s)$. His result was sharpened by Bollobás, Daykin and Erdős [1], they proved for $n\geq 2k^{3}s$. Subsequently, Hao, Loh and Sudakov [14] proved the EMC for $n\geq 3k^{2}s$.
Frankl [9] showed that $m(n,k,s)\leq{n\choose k}-{n-s\choose k}$ when $n\geq(2s+1)k-s$. Finally Frankl and Kupavskii prove the following theorem.
@@
In non-uniform case Kleitman considered when $s=3$.
@@
$|\mathcal{F}|\leq\sum_{j\leq m+1}^{3m+2}{3m+2\choose j}$
@@ 2. Results
The condition of Kleitman’s theorem can be written in the following form: if $|A\cap B|+|B\cap C|+|C\cap A|\geq 1$ holds for any 3 distinct members $A$, $B$, $C\in{\mathcal{F}}$. The main goal of the present paper is to study the case when 1 is replaced by $\ell$ on the right-hand side. Since the case $\ell=1$ has been studied a lot, we will only look at the case $\ell\geq 2$.
Let $d:(2^{[n]})^{3}\rightarrow\mathbb{N}$ be a function such that for all $A$, $B$, $C\in 2^{[n]}$, $d(A,B,C)=|A\cap B|+|B\cap C|+|C\cap A|$. Denote by $\mathbf{H}_{\ell}$ the class of families such that $\mathcal{H}\in\mathbf{H}_{\ell}$, $d(A,B,C)\geq\ell$ is true for all sets $A$, $B$, $C\in\mathcal{H}$. Let $g(n,k,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset{[n]\choose k},\mathcal{F}\in\mathbf{H_{\ell}}\}$ and $h(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\mathbf{H_{\ell}}\}$.
First we will consider the uniform case. Here we prove the following upper and lower bounds:
@@
$f(j,\ell)=\left\{\begin{array}[]{l l}2j-\lceil\frac{\ell}{3}\rceil&\text{if}\hskip 5.69054pt\frac{\ell}{3}\leq j<\frac{2\ell}{3},\vspace{1mm}\\ 3j-\ell&\text{if}\hskip 5.69054pt\frac{2\ell}{3}\leq j.\\ \end{array}\right.$
@@
If $a<b$, then we interpret (${a\choose b}$ as zero.
@@
It can be seen that this theorem is a kind of sharpening of the Erdős-Ko-Rado theorem, since it achieves the same upper bound for the corresponding pairs $(n,k)$ under weaker conditions.
In the non-uniform case, we investigate families $\mathcal{F}$ for which $d(A,B,C)\geq 3n-x$ for all $A$, $B$, $C\in\mathcal{F}$. In the case when $x=6p$ ($p$ is an integer) the following construction seems to be optimal: we select all sets with at least $n-p$ elements, then for any set of triples $d(A,B,C)\geq 3n-6p$. If $x=6p$, this will be the optimal solution for large $n$ and if $6p<x<6(p+1)$, then the optimal solution contains all $n-p$-element sets and some of the $n-p-1$-element sets.
@@ 3.1. Shifting a family
=$\mathcal{F}^{\prime}=\{F:F\in\mathcal{F},(x\notin F)\vee(x,y\in F)\vee(x\in F,y\notin F,F\cup\{y\}-\{x\}\in\mathcal{F})\}$
@@
′′=$\mathcal{F}^{\prime\prime}=\{F:F\in\mathcal{F},x\in F,y\notin F,F\cup\{y\}-\{x\}\notin\mathcal{F}\}.$
@@
Let ${\mathcal{F}}^{1}=\{H:H\in{\mathcal{F}},H=\tau(H)\}$ and ${\mathcal{F}}^{2}=\{H:H\in{\mathcal{F}},H\neq\tau(H)\}$. Let $\mathcal{A}^{1}=\mathcal{A}\cap{\mathcal{F}}^{1}$ and $\mathcal{A}^{2}=\mathcal{A}\cap{\mathcal{F}}^{2}$. We can assume that $\mathcal{A}^{1}$ contains the element $x$ $p$ times, the element $y$ $q$ times and $\mathcal{A}^{2}$ contains the element $x$ $r$ times. (It is obvious that $\mathcal{A}^{2}$ contain the element $y$ 00 times.) Then $d(\mathcal{A})=c+{p+r\choose 2}+{q\choose 2}>c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))$.
This is true only if $p+r>p>q$ (and $p+r>q+r>q$). This means that there are at least $p-q$ sets $H\in\mathcal{A}^{1}$ such that $x\in H$, $y\notin H$. Since $x\in\tau(H)$ and $y\notin\tau(H)$, then $H^{\prime}=H-\{x\}\cup\{y\}\in{\mathcal{F}}$. So if $\mathcal{A}^{\prime}$ is a family such that we replace $x$ with $y$ in this $p-q$ sets in $\mathcal{A}^{1}$, then $\mathcal{A}^{\prime}\subset{\mathcal{F}}$ and $d(\mathcal{A}^{\prime})=c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))<\ell$, a contradiction. ∎
@@
A uniform family ($\mathcal{F}\in{[n]\choose k}$ is called shifted, if $H\in\mathcal{F}$ and $I\leq H$ imply $I\in\mathcal{F}$. A non-uniform family $\mathcal{F}\in 2^{[n]}$ is called shifted, if ($\mathcal{F}\cap{[n]\choose i}$ is shifted for every $0\leq i\leq n$.
@@
As a result of Claim 3.4 and 3.10, if a family $\mathcal{F}\in\mathbf{H}_{\ell}$, then there exists a shifted family $\mathcal{F^{\prime}}\in\mathbf{H}_{\ell}$ such that $|\mathcal{F}|=|\mathcal{F}^{\prime}|$. Therefore it is enough to consider the shifted families, when we search for the maximum sized family in $\mathbf{H}_{\ell}$.
@@
Let $\ell=2$ or $3$, ($\mathcal{F}\subset{[n]\choose k}$ and $\mathcal{F}\in\mathbf{H}_{\ell}$ be a shifted family. If $A$, $B$, $C\in\mathcal{F}$ and $A^{\prime}=A\cap[3k-\ell]$, $B^{\prime}=B\cap[3k-\ell]$ and $C^{\prime}=C\cap[3k-\ell]$, then $d(A^{\prime},B^{\prime},C^{\prime})\geq\ell$, that is $\mathcal{F}\cap[3k-\ell]\in\mathcal{H}_{\ell}$.
@@
Claim 3.12 is not true if $\ell\geq 4$. For example let $\ell=4$, $k=3$, $A=\{1,2,6\}$, $B=\{1,3,6\}$, $C=\{1,4,5\}$ and $H\in\mathcal{F}$ if and only if $H\leq A$ or $H\leq B$ or $H\leq C$. Then $\mathcal{F}\in\mathbf{H}_{\ell}$, because for all $H\in{\mathcal{F}}$ are true that $1\in H$ and contains two element of $\{2,3,4,5,6\}$, so $d(\mathcal{A}\cap{1})=3$ and $d(\mathcal{A}-{1})\geq 1$, so $d(\mathcal{A})\geq 4$ but $53k-\ell=$ 5 and $d(A^{\prime},B^{\prime},C^{\prime})<4$.
@@ 3.3. Lower bound for the uniform case
Proof of Theorem 2.3. Let $j$ be an integer and let =$\mathcal{F}=\{F:F\in{[n]\choose k},|F\cap[f(j,\ell)]|\geq j\}$ be a family. Because of Claim 3.20 we know that $\mathcal{F}\in\mathbf{H}_{\ell}$. The expression (${f(j,\ell)\choose i}{n-f(j,\ell)\choose k-i}$ is the number of sets containing $i$ elements from $[3j-\ell]$. By choosing the optimal $j$ we get the lower bound. $\square$
@@
If $j=1$, then $f(j,\ell)=1$, so the value of the expression in (1) is ($n-1\choose k-1$. If $j=2$, then $f(j,\ell)=4$, so in this case the value of the expression in (1) is (${4\choose 2}{n-4\choose k-2}+{4\choose 3}{n-4\choose k-3}+{4\choose 4}{n-4\choose k-4}$. So $m(n,k,\ell)\geq 2$ if
@@
Proof of Theorem 2.5. First let us assume that there are two sets $A$ and $B\in\mathcal{F}$ that are disjoint. Since $|A\cap B|=0$, therefore $|A\cap C|+|B\cap C|\geq 2$, so all sets must contain at least 2 elements from $A\cup B$. An upper bound can be obtained by choosing the first two elements of the sets from $A\cup B$ and the others from $[n]$, so that ||($|\mathcal{F}|\leq{2k\choose 2}{n\choose k-2}$.
The other case when there are no two disjoint sets. In that case we can use Theorem 1.3, so ||($|\mathcal{F}|\leq{n-1\choose k-1}$ and equality is achieved if and only if there is an element $p\in[n]$ such that $H\in\mathcal{F}$ if and only if $p\in H$.
To prove the theorem we have to show that (${n-1\choose k-1}\geq{2k\choose 2}{n\choose k-2}$. Simplifying this, we get that $(n-k+2)(n-k+1)\geq nk(k-1)(2k-1)$. Using the condition $n\geq 4k^{3}$, we get that $(n-k+2)(n-k+1)\geq\frac{n^{2}}{2}\geq\frac{n}{2}4k^{3}\geq nk(k-1)(2k-1)$. $\square$

### v2

@@ Abstract.
We call a family $\mathcal{F}$ $(3,2,\ell)$-intersecting if $|A\cap B|+|B\cap C|+|C\cap A|\geq\ell$ for all $A$, $B$, $C\in\mathcal{F}$. We try to look for the maximum size of such a family $\mathcal{F}$ in case when $\mathcal{F}\subset{[n]\choose k}$ or $\mathcal{F}\subset 2^{[n]}$. In the uniform case we show that if $\mathcal{F}$ is $(3,2,2)$-intersecting, then $|\mathcal{F}|\leq{n+1\choose k-1}+{n\choose k-2}$ and if $\mathcal{F}$ is $(3,2,3)$-intersecting, then $|\mathcal{F}|\leq{n\choose k-1}+2{n\choose k-3}+3{n-1\choose k-3}$. For the lower bound we construct a $(3,2,\ell)$-intersecting family and we show that this bound is sharp when $\ell=2$ or $3$ and $n$ is sufficiently large compared to $k$. In the non-uniform case we give an upper bound for a (3,2,nx)(3,2,n-x)-intersecting family, when $n$ is sufficiently large compared to $x$.
@@ 1. Introduction
[2] Let $k$, $n\geq 2k$ be integers and ${\mathcal{F}}\subset{[n]\choose k}$ be an intersecting family. Then $|{\mathcal{F}}|\leq\displaystyle\binom{n-1}{k-1}$.
@@
It was proved by Erdős, Ko, and Rado in their seminal paper [2] that for any $1\leq t<k$, the largest $t$-intersecting families $\mathcal{F}\subset{[n]\choose k}$ are the maximal trivially $t$-intersecting ones if $n$ is large enough. The threshold function $n_{0}(k,t)=(k-t+1)(t+1)$ of this property was determined by Frankl [5] for $t\geq 15$. Later Wilson [19] filled in the gap $2\leq t\leq 14$.
@@
The non-uniform version of the theorem for $t\geq 2$ was proved by Katona.
@@
$|\mathcal{F}|\leq\left\{\begin{array}[]{c c}\sum_{i=\frac{n+t}{2}}^{n}{n\choose i}&\text{if $n+t$ is even}\vskip 2.84526pt\\ \vskip 2.84526pt\sum_{i=\frac{n+t+1}{2}}^{n}{n\choose i}+{n-1\choose\frac{n+t-1}{2}}&\text{if $n+t$ is odd}\end{array}\right$.
@@
[3] For $n\geq k(s+1)$,
$m(n,k,s)=\max\left\{{n\choose k}-{n-s\choose k},{k(s+1)-1\choose k}\right\}$.
It was one of the favourite problems of Erdős and there was hardly a combinatorial lecture of him where he did not mention it. The EMC is trivial for $k=1$ and was proved by Erdős and Gallai [4] for $k=2$ then it was settled in the case $k=3$ and $n\geq 4s$ in [13], for $k=3$, all $n$ and $s\geq s_{0}$ in [18], and finally, it was completely resolved for $k=3$ in [8].
The case $s=1$ is the classical Erdős-Ko-Rado theorem. In his original paper, Erdős proved 1.8 for $n\geq n_{0}(k,s)$ for some $n_{0}(k,s)$. His result was sharpened by Bollobás, Daykin and Erdős [1], they proved for $n\geq 2k^{3}s$. Subsequently, Hao, Loh and Sudakov [15] proved the EMC for $n\geq 3k^{2}s$.
Frankl [9] showed that $m(n,k,s)\leq{n\choose k}-{n-s\choose k}$ when $n\geq(2s+1)k-s$. For $s$ very large Frankl and Kupavskii prove the following theorem.
@@
In the non-uniform case Kleitman considered when $s=2$.
@@
$|\mathcal{F}|\leq\sum_{j=m+1}^{3m+2}{3m+2\choose j}$
is the best upper bound.
Theorem 1.11.
[17] Let $n=3m$ and $\mathcal{F}\subset 2^{[n]}$ be a family with no three of whose members are pairwise disjoint. Then
$|\mathcal{F}|\leq\sum_{j=m+1}^{3m}{3m\choose j}+\frac{2}{3}{3m\choose m}$
@@ 2. Results
The condition of Kleitman’s theorem can be written in the following form: $|A\cap B|+|B\cap C|+|C\cap A|\geq 1$ holds for any 3 distinct members $A$, $B$, $C\in{\mathcal{F}}$. The main goal of the present paper is to study the case when 1 is replaced by $\ell$. Since the case $\ell=1$ has been studied a lot, we will only look at the case $\ell\geq 2$.
Let $d:(2^{[n]})^{3}\rightarrow\mathbb{N}$ be a function such that for all $A$, $B$, $C\in 2^{[n]}$, $d(A,B,C)=|A\cap B|+|B\cap C|+|C\cap A|$. Denote by $\mathbf{H}_{\ell}$ the class of families such that $\mathcal{H}$, $d(A,B,C)\geq\ell$ for all sets $A$, $B$, $C\in\mathcal{H}$. Let $g(n,k,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset{[n]\choose k},\mathcal{F}\in\mathbf{H_{\ell}}\}$ and $h(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\mathbf{H_{\ell}}\}$.
First we consider the uniform case. We prove the following upper and lower bounds:
@@
[…]
@@
If $a<b$, then we interpret $a\choose b$} as zero.
@@
This theorem is a sharpening of the Erdős-Ko-Rado theorem, since it achieves the same upper bound for the corresponding pairs $(n,k)$ under weaker conditions.
In the non-uniform case, we investigate families $\mathcal{F}$ for which $d(A,B,C)\geq 3n-x$ for all $A$, $B$, $C\in\mathcal{F}$. In the case when $x=6p$ ($p$ is an integer) the following construction seems to be optimal: we select all sets with at least npn-p elements, then for any set of triples $d(A,B,C)\geq 3n-6p$. If $x=6p$, this will be the optimal solution for large $n$ and if $6p<x<6(p+1)$, then the optimal solution contains all npn-p-element sets and some of the np1n-p-1-element sets.
@@ 3.1. Shifting a family
$\mathcal{F}^{\prime}=\{F:F\in\mathcal{F},(x\notin F)\vee(x,y\in F)\vee(x\in F,y\notin F,F\cup\{y\}-\{x\}\in\mathcal{F})\}$
@@
$\mathcal{F}^{\prime\prime}=\{F:F\in\mathcal{F},x\in F,y\notin F,F\cup\{y\}-\{x\}\notin\mathcal{F}\}.$
@@
Let ${\mathcal{F}}^{1}=\{H:H\in{\mathcal{F}},H=\tau(H)\}$ and ${\mathcal{F}}^{2}=\{H:H\in{\mathcal{F}},H\neq\tau(H)\}$. Let $\mathcal{A}^{1}=\mathcal{A}\cap{\mathcal{F}}^{1}$ and $\mathcal{A}^{2}=\mathcal{A}\cap{\mathcal{F}}^{2}$. We can assume that $\mathcal{A}^{1}$ contains the element $x$ $p$ times, the element $y$ $q$ times and $\mathcal{A}^{2}$ contains the element $x$ $r$ times. (It is obvious that $\mathcal{A}^{2}$ contain the element $y$ 0 times.) Then $d(\mathcal{A})=c+{p+r\choose 2}+{q\choose 2}>c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))$.
This is true only if $p+r>p>q$ (and $p+r>q+r>q$). This means that there are at least pqp-q sets $H\in\mathcal{A}^{1}$ such that $x\in H$, $y\notin H$. Since $x\in\tau(H)$ and $y\notin\tau(H)$, then $H^{\prime}=H-\{x\}\cup\{y\}\in{\mathcal{F}}$. So if $\mathcal{A}^{\prime}$ is a family such that we replace $x$ with $y$ in this pqp-q sets in $\mathcal{A}^{1}$, then $\mathcal{A}^{\prime}\subset{\mathcal{F}}$ and $d(\mathcal{A}^{\prime})=c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))<\ell$, a contradiction. ∎
@@
A uniform family $\mathcal{F}\in{[n]\choose k}$ is called shifted, if $H\in\mathcal{F}$ and $I\leq H$ imply $I\in\mathcal{F}$. A non-uniform family $\mathcal{F}\in 2^{[n]}$ is called shifted, if $\mathcal{F}\cap{[n]\choose i}$ is shifted for every $0\leq i\leq n$.
@@
As a result of Claim 3.4 and 3.10, if a family $\mathcal{F}\in\mathbf{H}_{\ell}$, then there exists a shifted family $\mathcal{F^{\prime}}\in\mathbf{H}_{\ell}$ such that $|\mathcal{F}|=|\mathcal{F}^{\prime}|$. Therefore it is sufficient to consider the shifted families, when we search for the maximum sized family in $\mathbf{H}_{\ell}$.
@@
Let $\ell=2$ or $3$, $\mathcal{F}\subset{[n]\choose k}$ and $\mathcal{F}\in\mathbf{H}_{\ell}$ be a shifted family. If $A$, $B$, $C\in\mathcal{F}$ and $A^{\prime}=A\cap[3k-\ell]$, $B^{\prime}=B\cap[3k-\ell]$ and $C^{\prime}=C\cap[3k-\ell]$, then $d(A^{\prime},B^{\prime},C^{\prime})\geq\ell$, that is $\mathcal{F}\cap[3k-\ell]\in\mathcal{H}_{\ell}$.
@@
Claim 3.12 is not true if $\ell\geq 4$. For example let $\ell=4$, $k=3$, $A=\{1,2,6\}$, $B=\{1,3,6\}$, $C=\{1,4,5\}$ and $H\in\mathcal{F}$ if and only if $H\leq A$ or $H\leq B$ or $H\leq C$. Then $\mathcal{F}\in\mathbf{H}_{\ell}$, because for all $H\in{\mathcal{F}}$ are true that $1\in H$ and contains two element of $\{2,3,4,5,6\}$, so $d(\mathcal{A}\cap{1})=3$ and $d(\mathcal{A}-{1})\geq 1$, so $d(\mathcal{A})\geq 4$ but $3k-\ell=5$ and $d(A^{\prime},B^{\prime},C^{\prime})<4$.
@@ 3.3. Lower bound for the uniform case
Proof of Theorem 2.3. Let $j$ be an integer and let $\mathcal{F}=\{F:F\in{[n]\choose k},|F\cap[f(j,\ell)]|\geq j\}$ be a family. Because of Claim 3.20 we know that $\mathcal{F}\in\mathbf{H}_{\ell}$. The expression ${f(j,\ell)\choose i}{n-f(j,\ell)\choose k-i}$ is the number of sets containing $i$ elements from $3j-\ell]$. By choosing the optimal $j$ we get the lower bound. $\square$
@@
If $j=1$, then $f(j,\ell)=1$, so the value of the expression in (1) is $n-1\choose k-1$. If $j=2$, then $f(j,\ell)=4$, so in this case the value of the expression in (1) is ${4\choose 2}{n-4\choose k-2}+{4\choose 3}{n-4\choose k-3}+{4\choose 4}{n-4\choose k-4}$. So $m(n,k,\ell)\geq 2$ if
@@
Proof of Theorem 2.5. First let us assume that there are two sets $A$ and $B\in\mathcal{F}$ that are disjoint. Since $|A\cap B|=0$, therefore $|A\cap C|+|B\cap C|\geq 2$, so all sets must contain at least 2 elements from $A\cup B$. An upper bound can be obtained by choosing the first two elements of the sets from $A\cup B$ and the others from $[n]$, so that $|\mathcal{F}|\leq{2k\choose 2}{n\choose k-2}$.
The other case when there are no two disjoint sets. In that case we can use Theorem 1.3, so $|\mathcal{F}|\leq{n-1\choose k-1}$ and equality is achieved if and only if there is an element $p\in[n]$ such that $H\in\mathcal{F}$ if and only if $p\in H$.
To prove the theorem we have to show that ${n-1\choose k-1}\geq{2k\choose 2}{n\choose k-2}$. Simplifying this, we get that $(n-k+2)(n-k+1)\geq nk(k-1)(2k-1)$. Using the condition $n\geq 4k^{3}$, we get that $(n-k+2)(n-k+1)\geq\frac{n^{2}}{2}\geq\frac{n}{2}4k^{3}\geq nk(k-1)(2k-1)$. $\square$